OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 + )12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $2,000 at an interest rate of 10% for 4 years, while Jenny invests $2,000 at an interest rate of 5% for 8 years.

OpenStudy (anonymous):

determine the amount gained by jenny and jimmy

OpenStudy (anonymous):

something's missing in your formula...

OpenStudy (anonymous):

its P(1+r/12)^12t

hero (hero):

The general formula for compound interest: \[A = P(1 + \frac{r}{n})^{nt}\]

hero (hero):

For Jimmy: P = 2000 n = 12 r = .10 t = 4 For Jenny P = 2000 n = 12 r = .05 t = 8 r = interest rate n = number of months in a year interest is compounded t = number of years of investment

hero (hero):

You should be able to figure it out from there.

OpenStudy (anonymous):

how did you arrive at those answers?

hero (hero):

From the given information in the problem and also from knowing the general formula of compound interest and what each variable represents

OpenStudy (anonymous):

oh im still lost but its ok/

hero (hero):

All you do is replace the variables with the appropriate numbers then compute A via calculator

hero (hero):

For Jimmy, After substitution: \[A = 2000(1+ \frac{.10}{12})^{12 \times 4}\]